# HARMONIOUS – Human-like reactive motion control and multimodal perception for humanoid robots

### v1

@@ II-A Experimental setup
Figure 2: HARMONIOUS – overview. The blue box represents the local trajectory sampling. The trajectory towards the target pose is sampled using a minimum-jerk generator (position) and a low-pass filter (orientation). The target pose is received from an external target generator. The green box represents the obstacle processing. It takes inputs from proximity sensors, skin parts (tactile stimuli), and RGB-D camera (visual stimuli) and computes linear inequality constraints to dodge obstacles or react to the collision that occurred. The obstacle exists virtually for a specific survival time with decreasing threat level, giving humans time to react. The red box represents the controller with a QP solver that solves the differential kinematics problem. It takes the next desired pose and current joint positions as input. If the problem is feasible, the computed joint velocities are integrated into new joint positions and sent to the robot. Otherwise, the position constraint is relaxed, and the QP problem is solved again.
II-B Motion controller
The main component of our approach is a motion controller (see the red box in Fig. 2). Our motion controller is based on a reactive kinematic control, in particular the approach called differential kinematics which is defined as
@@ II-B Motion controller
where $\boldsymbol{\nu}$ is the spatial velocity of the end effector, $\dot{\mathbf{q}}$ are the joint velocities of the robot and $\mathbf{J}(\mathbf{q})$ is the Jacobian of the robot for the joint positions $\mathbf{q}$. This relation represents a trajectory constraint between robot configuration and Cartesian space.
@@
$\boldsymbol{\nu^{pos}}=\frac{\mathbf{x_{t}}-\mathbf{x_{c}}}{t_{s}},\qquad\boldsymbol{\nu^{ori}}=\frac{\mathbf{R_{t}}\mathbf{R_{c}}^{T}}{t_{s}}$ (2)
where $\boldsymbol{\nu^{pos}}$ is the translational velocity of the end effector, $\mathbf{x_{c}}$ and $\mathbf{x_{t}}$ are the current and desired next end effector position, $\boldsymbol{\nu^{ori}}$ is the rotational velocity of the end effector, $\mathbf{R_{c}}$ and $\mathbf{R_{t}}$ are the current and desired next end effector orientation, and $t_{s}$ is the period time.
For our purpose, we can relax the trajectory constraint by adding a slack variable $\boldsymbol{\lambda}$ to help the controller with finding a solution when other (e.g., safety) constraints appear:
@@
Then, we formulate the problem as a strictly convex QP problem with joint velocities $\dot{\mathbf{q}}$ and slack variables $\boldsymbol{\lambda}$ as optimization variables. In our cost function, we minimize the sum of the squared Euclidean norm of the joint velocities, the distance from the natural joint positions, and the squared Euclidean norm of the vector of slack variables. It is formulated as:
$\min_{\dot{\mathbf{q}},\boldsymbol{\lambda}}\ \frac{\mu}{2}\dot{\mathbf{q}}^{T}\mathbf{W_{q}}\dot{\mathbf{q}}+\frac{1}{2}\boldsymbol{\lambda}^{T}\mathbf{W_{\lambda}}\boldsymbol{\lambda}+\frac{c_{h}}{2}\left(\dot{\mathbf{q}}-\dot{\mathbf{q}}_{n}\right)^{T}\mathbf{W_{q}}\left(\dot{\mathbf{q}}-\dot{\mathbf{q}}_{h}\right)$ (4)
where diagonal matrices $\mathbf{W_{q}}$ and $\mathbf{W_{\lambda}}$ contain the weights of the individual joints and the relaxations for the trajectory constrain, $\mathbf{\dot{q}}_{h}$ are the velocities needed to reach natural (home) robot configuration, $c_{h}$ is a weight of the motivation to keep robot close to his natural configuration, and $\mu$ is a damping factor calculated from the manipulability index $\omega=\sqrt{\mathbf{JJ}^{T}}$ [49] as:
@@
This damping factor increases when the manipulability index decreases to zero, implying that the robot is close to a singularity position (the manipulability index is zero in that position). The higher damping factor makes the minimization of joint velocities more important, leading to lower joint velocities and singularity avoidance [27].
Apart from the mentioned relaxed trajectory constrained, the minimization is done with respect to several other constraints, such as joint velocity and position bounds (see Sec. II-K), slack variable bounds, collision avoidance (see Sec. II-F), and other task or kinematic constraints (see Sec. II-J). The constraints can be written as follows:
$\mathrm{s.t.}\quad\begin{cases}\boldsymbol{\nu}-\boldsymbol{\lambda}=\mathbf{J}(\mathbf{q})\dot{\mathbf{q}}\\ \dot{\mathbf{q}}_{L}<\dot{\mathbf{q}}<\dot{\mathbf{q}}_{U}\\ \mathbf{q}_{L}<\mathbf{q}<\mathbf{q}_{U}\\ -\boldsymbol{\varepsilon}<\boldsymbol{\lambda}<\boldsymbol{\varepsilon}\\ \mathbf{A_{o}}\leq\mathbf{b_{o}}\\ \mathbf{b_{L,t}}\leq\mathbf{A_{t}}\leq\mathbf{b_{U,t}}\end{cases}$ (6)
where $\mathbf{\dot{q}}_{L,U},\mathbf{q}_{L,U},\boldsymbol{\varepsilon}$ are the lower and upper bounds for the joint velocities, the joint positions, and the slack variables, respectively. $\mathbf{A_{o}}$ and $\mathbf{b_{o}}$ define obstacles avoidance constraints; $\mathbf{A_{t}}$, $\mathbf{b}_{L,t}$, and $\mathbf{b}_{U,t}$ define other task constraints.
II-C Dual-arm trajectory constraint
Here, we formulate the problem for two arms with a common torso. Since we have a common torso of the robot, we cannot calculate the trajectory constraints for the arms independently of each other. Moreover, it is not always possible to reach targets with both arms simultaneously, conflict situations are managed by prioritizing one of the arms and calling it the “primary arm”, and the other is then the “secondary arm”. Therefore, the trajectory constraint (Eq. 3) for a robot with a common torso is expanded as:
@@ II-C Dual-arm trajectory constraint
where the subscripts $t,p,s$ represent torso, “primary arm”, and “secondary arm” respectively. The superscripts $pos,ori$ represent the relaxation of the constraint of the position and orientation trajectory, respectively. The single-arm formulation contains only the torso and the “primary arm” part. $\mathbf{J^{p/s}_{a,pos}}(\mathbf{q^{p/s}}),\mathbf{J^{p/s}_{t,pos}}(\mathbf{q^{t}})$ are the translational velocity components of the Jacobian matrices for arm and torso, respectively, $\mathbf{J^{p/s}_{a,ori}}(\mathbf{q^{p/s}}),\mathbf{J^{p/s}_{t,ori}}(\mathbf{q^{t}})$ are the rotational velocity components of the Jacobian matrices for arm and torso, respectively.
Consistent with the arm prioritization, we distinguish two cases for the limits of the slack variable $\boldsymbol{\varepsilon}$. They are set to $\infty$ (the deviation is minimized only) or to zero (the equality constraint without relaxation). In our solution, $\boldsymbol{\varepsilon}$ is set to $\infty$ for all $\boldsymbol{\lambda}$ except $\boldsymbol{\lambda_{p}^{pos}}=[0,0,0]^{T}$. If the equality constraint on the desired translation velocity of the “primary arm” makes the problem infeasible, we solve the problem again with the constraint relaxed to the minimization of the deviation. This strategy allows the robot to reach a target position with higher precision if the original problem is feasible and to make it feasible if the equality constraint is too strict.
II-D Local trajectory sampling
@@ II-D1 Position sampling
where $a\approx-150.766$, $b\approx-84.981$, and $c\approx-15.967$ are the constant coefficients found by the optimization, and $T$ is a parameter (related to the execution time) that regulates the reactivity of the filter. Throughout our experiments, we set $T=\left(||\mathbf{x_{d}}-\mathbf{x_{0}}||/v_{t}\right)$ s, where $\mathbf{x_{d}}$ is the desired target position, $\mathbf{x_{0}}$ is the starting position, and $v_{t}$ is the desired Cartesian speed of movement.
The output $\mathbf{x}\left(t\right)$ of the filter in Eq. 9 provides a sampling of the input $\mathbf{x}_{d}\left(t\right)$ that is “quasi” minimum-jerk, being the filter only an approximation of the system (Eq. 8), although we proved that it is reliable and effective in smoothing out sharp transitions while guaranteeing bell-shaped velocity profiles.
II-D2 Orientation sampling
We use the axis-angle representation for the orientations of the end effector. The orientation is described by a rotation vector $\mathbf{r}=\left[r_{1},r_{2},r_{3}\right]^{T},$ that encodes the axis of rotation $u=\mathbf{r}/||\mathbf{r}||$ and a rotation angle $\Theta=||\mathbf{r}||$. Orientation sampling is carried out by interpolation between two orientations $\mathbf{r_{1}},\mathbf{r_{2}}$ using a spherical linear interpolation, i.e., Slerp [52], which takes advantage of the fact that the spherical metric of $S^{3}$ is the same as the angular metric of $SO(3)$ for a constant angular speed of movement [52]. Moreover, it creates a torque-minimal path for the rotation as it travels along the straightest path the rounded surface of a sphere and it.
We first start by creating skew matrices $\mathbf{r_{+,1}},\mathbf{r_{+,2}}$ from rotation vectors $\mathbf{r_{1}},\mathbf{r_{2}}$ as
$\mathbf{r_{+}}=\begin{bmatrix}0&-r_{3}&r_{2}\cr r_{3}&0&-r_{1}\cr-r_{2}&r_{1}&0\end{bmatrix}.$
These matrices can be easily converted to rotation matrices $\mathbf{R_{1}},\mathbf{R_{2}}$ by computing their matrix exponentials (expm function):
@@ II-D2 Orientation sampling
We compute the interpolation between rotations $\mathbf{R_{1}},\mathbf{R_{2}}$ as
@@
II-E Peripersonal space projection
@@
II-F Obstacle processing
The processed obstacles are the second input to the reactive controller (see the green box in Fig. 2). In our setup, we operate with obstacles detected by visual, proximity, and tactile sensors. However, our approach is universal in modalities. The processed obstacles, called collision points ($C$), are represented by their projected position on the robot $\mathbf{P}_{C}$, collision direction $\mathbf{n}_{C}$, threat level $a_{t}$ (i.e., scaled sensor measurement) and a gain $V_{C}$ representing the severity of the information; for example, tactile events can have the highest gain as the collision has already happened. The main advantage of our representation is that it jointly enables the post-collision reaction (tactile sensors) and collision avoidance (proximity and visual sensors).
An obstacle is incorporated into the QP problem as a linear inequality constraint that restricts movements towards the obstacle but allows motion in tangential directions. The constraint is defined as a limitation of the approach motion of the collision point $C$ with Cartesian velocity $\mathbf{\dot{x}}_{C}$ towards the obstacle (in direction $\mathbf{n}_{C}$) by a maximum approach velocity $\mathbf{\dot{x}_{a}}$ as
$\mathbf{n}^{T}_{C}\mathbf{\dot{x}_{C}}=\mathbf{n}_{C}^{T}\mathbf{J}_{C}\mathbf{\dot{q}_{C}}<\mathbf{\dot{x}_{a}}$ (11)
where $\mathbf{J}_{C}$ is the translational-velocity component of the Jacobian associated with $C$ and $\mathbf{\dot{q}}_{C}$ are the joint velocities of the kinematic chain ending with $C$. The maximum approach velocity $\mathbf{\dot{x}_{a}}$ is calculated from the obstacle threat level $a_{t}$ as
$\mathbf{\dot{x}_{a}}=(k_{1}-V_{C}\cdot a_{t})\cdot k_{2}$ (12)
where $k_{1}$ and $k_{2}$ are positive values to adjust the constraint. In accordance with Eq. 4, we can write the obstacle constraints in the matrix form ($\mathbf{A^{i}_{o}}\leq b^{i}_{o}$) as

### v2

@@ III-A Experimental setup
Figure 2: HARMONIOUS – overview. The blue box represents the local trajectory sampling. The trajectory towards the target pose is sampled using a minimum-jerk generator (position) and a low-pass filter (orientation). The target pose is received from an external target generator. The green box represents the obstacle processing. It takes inputs from proximity sensors, skin parts (tactile stimuli), and RGB-D camera (visual stimuli) and computes linear inequality constraints to dodge obstacles or react to the collision that occurred. The obstacle exists virtually for a specific surviving time with a decreasing threat level, giving humans time to react. The red box represents the controller with a QP solver that solves the differential kinematics problem. It takes the next desired pose and current joint positions as input. If the problem is feasible, the computed joint velocities are integrated into new joint positions and sent to the robot. Otherwise, the position constraint is relaxed, and the QP problem is solved again.
III-B Motion controller
The main component of our approach is a motion controller (see the red box in Fig. 2). Our motion controller is based on reactive kinematic control, in particular the approach called differential kinematics which is defined as
@@ III-B Motion controller
where $\boldsymbol{\nu}\in\mathbb{R}^{m}$ is the spatial velocity of the end effector ($m$ is the task dimension—6 for single-arm version, 12 for dual-arm version), $\dot{\mathbf{q}}\in\mathbb{R}^{n}$ are the joint velocities of the robot ($n$ is the number of joints) and $\mathbf{J}(\mathbf{q})\in\mathbb{R}^{m\times n}$ is the Jacobian of the robot for the joint positions $\mathbf{q}\in\mathbb{R}^{n}$. This relation represents a trajectory constraint between the robot configuration and Cartesian space. In this work, we assume the reference frame for Cartesian coordinates and Jacobians is the base frame of the robot (around the waist).
@@
$\boldsymbol{\nu^{pos}}=\frac{\mathbf{x_{t}}-\mathbf{x_{c}}}{t_{s}},\qquad\boldsymbol{\nu^{ori}}=\frac{\varphi(\mathbf{R_{t}}\mathbf{R_{c}}^{T})}{t_{s}}$ (2)
where $\boldsymbol{\nu^{pos}}\in\mathbb{R}^{3}$ is the translational velocity of the end effector, $\mathbf{x_{c}}\in\mathbb{R}^{3}$ and $\mathbf{x_{t}}\in\mathbb{R}^{3}$ are the current and desired next end effector position, $\boldsymbol{\nu^{ori}}\in\mathbb{R}^{3}$ is the rotational velocity of the end effector, $\mathbf{R_{c}}\in\mathbb{R}^{3\times 3}$ and $\mathbf{R_{t}}\in\mathbb{R}^{3\times 3}$ are the current and desired next end effector orientation (converted from axis-angle representation), $\varphi$ represents mapping from a rotation matrix to an axis-angle representation to obtain angular velocities using a period time $t_{s}$ exploiting the infinitesimal rotation assumption (see e.g., [2]).
For our purposes, we can relax the trajectory constraint (the “reaching task”) by adding a slack variable $\boldsymbol{\lambda}\in\mathbb{R}^{m}$ to help the controller with finding a solution when other (e.g., safety) constraints appear:
@@
Then, we formulate the problem as a strictly convex QP problem with joint velocities $\dot{\mathbf{q}}$ and slack variables $\boldsymbol{\lambda}$ as optimization variables. In our cost function, we minimize the sum of the squared Euclidean norm of the joint velocities, the distance from the preferred joint positions, and the squared Euclidean norm of the vector of slack variables. It is formulated as:
$\min_{\dot{\mathbf{q}},\boldsymbol{\lambda}}\ \frac{\mu}{2}\dot{\mathbf{q}}^{T}\mathbf{W_{q}}\dot{\mathbf{q}}+\frac{1}{2}\boldsymbol{\lambda}^{T}\mathbf{W_{\lambda}}\boldsymbol{\lambda}+\frac{c_{h}}{2}\left(\dot{\mathbf{q}}-\dot{\mathbf{q}}_{n}\right)^{T}\mathbf{W_{q}}\left(\dot{\mathbf{q}}-\dot{\mathbf{q}}_{n}\right)$ (4)
where diagonal matrices $\mathbf{W_{q}}\in\mathbb{R}^{n\times n}$ and $\mathbf{W_{\lambda}}\in\mathbb{R}^{m\times m}$ contain the weights of the individual joints and the relaxations for the trajectory constraint, $\mathbf{\dot{q}}_{n}$ are the velocities needed to reach the preferred (e.g., home) robot configuration, $c_{h}$ is a weight of the motivation to keep robot close to a preferred configuration, and $\mu$ is a damping factor calculated as:
@@
where $\omega$ is a manipulability index [56] computed as $\omega=\sqrt{\mathrm{det}(\mathbf{J}(\mathbf{q})\mathbf{J(\mathbf{q})}^{T})}=s_{1}s_{2}...s_{n}$, where $s_{i}$ is the i-th singular value of $\mathbf{J}(\mathbf{q})$. This damping factor increases when the manipulability index decreases to zero, implying that the robot is close to a singularity position (the manipulability index is zero in that position). The higher damping factor makes the minimization of joint velocities more important, leading to lower joint velocities and singularity avoidance [29]. We assume that the weights ($\mathbf{W_{q}}$ $\mathbf{W_{\lambda}}$, $c_{h}$, and $\mu$) are non-negative to ensure that the problem is convex. The weights in $\mathbf{W_{q}}$ corresponding to the torso joints are higher than for the arm joints, which, through the minimization term, motivates smaller torso movements, which is more natural.
Apart from the mentioned relaxed trajectory constraint, the minimization is done with respect to several other constraints, such as joint velocity and position bounds (see Sec. III-K), slack variable bounds, collision avoidance (see Sec. III-F), and other task or kinematic constraints (see Sec. III-J). The constraints can be written as follows:
$\mathrm{s.t.}\quad\begin{cases}\boldsymbol{\nu}-\boldsymbol{\lambda}=\mathbf{J}(\mathbf{q})\dot{\mathbf{q}}\\ \dot{\mathbf{q}}_{L}<\dot{\mathbf{q}}<\dot{\mathbf{q}}_{U}\\ \mathbf{q}_{L}<\mathbf{q}<\mathbf{q}_{U}\\ -\boldsymbol{\varepsilon}<\boldsymbol{\lambda}<\boldsymbol{\varepsilon}\\ \mathbf{A_{o}}\mathbf{\dot{q}}\leq\mathbf{b_{o}}\\ \mathbf{b_{L,t}}\leq\mathbf{A_{t}}\mathbf{\dot{q}}\leq\mathbf{b_{U,t}}\end{cases}$ (6)
where $\mathbf{\dot{q}}_{L,U}\in\mathbb{R}^{n},\mathbf{q}_{L,U}\in\mathbb{R}^{n},\boldsymbol{\varepsilon}\in\mathbb{R}^{m}$ are the lower and upper bounds for the joint velocities, the joint positions, and the slack variables, respectively. $\mathbf{A_{o}}\in\mathbb{R}^{m_{o}\times n}$ and $\mathbf{b_{o}}\in\mathbb{R}^{m_{o}}$ define ($m_{o}$) obstacle avoidance constraints; $\mathbf{A_{t}}\in\mathbb{R}^{m_{t}\times n}$, $\mathbf{b}_{L,t}\in\mathbb{R}^{m_{t}}$, and $\mathbf{b}_{U,t}\in\mathbb{R}^{m_{t}}$ define other ($m_{t}$) task constraints.
III-C Dual-arm trajectory constraint
Here, we formulate the problem for two arms with a common torso. Since we have a common torso of the robot, we cannot calculate the trajectory constraints for the arms independently of each other. Moreover, it is not always possible to reach targets with both arms simultaneously. Conflict situations are managed by prioritizing one of the arms and calling it the “primary arm”, and the other is then the “secondary arm”. Therefore, the trajectory constraint (Eq. 3) for a robot with a common torso is expanded as:
@@ III-C Dual-arm trajectory constraint
where the subscripts $t,p,s$ represent torso, “primary arm”, and “secondary arm” respectively. The superscripts $pos,ori$ represent the relaxation of the constraint of the position and orientation trajectory, respectively. The single-arm formulation contains only the torso and the “primary arm” part. $\mathbf{J^{p/s}_{a,pos}}(\mathbf{q^{p/s}})\in\mathbb{R}^{3\times(n-3)},\mathbf{J^{p/s}_{t,pos}}(\mathbf{q^{t}})\in\mathbb{R}^{3\times 3}$ are the translational velocity components of the Jacobian matrices for arm and torso, respectively, $\mathbf{J^{p/s}_{a,ori}}(\mathbf{q^{p/s}})\in\mathbb{R}^{3\times(n-3)},\mathbf{J^{p/s}_{t,ori}}(\mathbf{q^{t}})\in\mathbb{R}^{3\times 3}$ are the rotational velocity components of the Jacobian matrices for arm and torso, respectively.
Consistent with the arm prioritization, we distinguish two cases for the limits of the slack variable $\boldsymbol{\varepsilon}$. They are set to $\infty$ (the deviation from the trajectory is minimized only) or to zero (the equality constraint without relaxation). In our solution, $\boldsymbol{\varepsilon}$ is set to $\infty$ for all $\boldsymbol{\lambda}$ except $\boldsymbol{\lambda_{p}^{pos}}=[0,0,0]^{T}$. If the equality constraint on the desired translation velocity of the “primary arm” makes the problem infeasible, we solve the problem again with the constraint relaxed to the minimization of the deviation. This strategy allows the robot to reach a target position with higher precision if the original problem is feasible and to make it feasible if the equality constraint is too strict.
III-D Local trajectory sampling
@@ III-D1 Position sampling
where $a\approx-150.766$, $b\approx-84.981$, and $c\approx-15.967$ are the constant coefficients found by the optimization, and $T$ is a parameter (related to the execution time) that regulates the reactivity of the filter. Throughout our experiments, we set $T=\left(||\mathbf{x_{d}}-\mathbf{x_{0}}||/v_{t}\right)$ s, where $\mathbf{x_{0}}\in\mathbb{R}^{3}$ is the starting position, and $v_{t}$ is the desired Cartesian speed of movement.
The output $\mathbf{x}\left(t\right)\in\mathbb{R}^{3}$ of the filter in Eq. 9 provides a sampling of the input $\mathbf{x}_{d}\left(t\right)$ that is “quasi” minimum-jerk, being the filter only an approximation of the system (Eq. 8), although the authors in [8] proved that it is reliable and effective in smoothing out sharp transitions while guaranteeing bell-shaped velocity profiles.
III-D2 Orientation sampling
We use the axis-angle representation for the orientations of the end effector. The orientation is described by a rotation vector $\mathbf{r}=\left[r_{1},r_{2},r_{3}\right]^{T},$ that encodes the axis of rotation $u=\mathbf{r}/||\mathbf{r}||$ and a rotation angle $\Theta=||\mathbf{r}||$. Orientation sampling is carried out by interpolation between two orientations $\mathbf{r_{1}},\mathbf{r_{2}}$ using a spherical linear interpolation, i.e., Slerp [59], which takes advantage of the fact that the spherical metric of $S^{3}$ is the same as the angular metric of $SO(3)$ for a constant angular speed of movement [59]. Moreover, it travels along the straightest (and shortest) path on the rounded surface of the quaternion unit sphere.
We first start by creating skew matrices $\mathbf{r_{+,1}},\mathbf{r_{+,2}}$ from rotation vectors $\mathbf{r_{1}},\mathbf{r_{2}}$. These matrices can be easily converted to rotation matrices $\mathbf{R_{1}},\mathbf{R_{2}}$ by computing their matrix exponentials (expm function):
@@ III-D2 Orientation sampling
We compute the interpolation between rotations $\mathbf{R_{1}},\mathbf{R_{2}}\in\mathbb{R}^{3\times 3}$ as
@@
III-E Peripersonal space projection
@@
III-F Obstacle processing
The processed obstacles are the second input to the reactive controller (see the green box in Fig. 2). In our setup, we operate with obstacles detected by visual, proximity, and tactile sensors. However, our approach is universal in modalities. The processed obstacles, called collision points ($C$), are represented by their projected position on the robot $\mathbf{P}_{C}\in\mathbb{R}^{3}$, collision direction $\mathbf{n}_{C}\in\mathbb{R}^{3}$, threat level $a_{t}$ (i.e., scaled sensor measurement) and a gain $V_{C}$ representing the severity of the information; for example, tactile events can have the highest gain as the collision has already happened. The main advantage of our representation is that it jointly enables the post-collision reaction (tactile sensors) and collision avoidance (proximity and visual sensors).
An obstacle is incorporated into the QP problem as a linear inequality constraint that restricts movements towards the obstacle but allows motion in tangential directions. The constraint limits the approach motion of the collision point $C$ with Cartesian velocity $\mathbf{\dot{x}}_{C}\in\mathbb{R}^{3}$ towards the obstacle (in direction $\mathbf{n}_{C}$) by a maximum approach velocity $\dot{x}_{a}$ (which can be negative and become the repulsive velocity, Sec. III-G) as
$\mathbf{n}^{T}_{C}\mathbf{\dot{x}_{C}}=\mathbf{n}_{C}^{T}\mathbf{J}_{C}\mathbf{\dot{q}_{C}}<\dot{x}_{a}$ (11)
where $\mathbf{J}_{C}\in\mathbb{R}^{3\times n_{c}}$ is the translational-velocity component of the Jacobian associated with $C$ and $\mathbf{\dot{q}}_{C}\in\mathbb{R}^{n_{c}}$ are the joint velocities of the kinematic chain ending with $C$. The maximum approach velocity $\dot{x}_{a}$ is calculated from the obstacle threat level $a_{t}$ as
$\dot{x}_{a}=(k_{1}-V_{C}\cdot a_{t})\cdot k_{2}$ (12)
where $k_{1}$ and $k_{2}$ are positive values to adjust the constraint which are described in Sec. III-G. In accordance with Eq. 4, we can write the obstacle constraints in matrix form ($\mathbf{A^{i}_{o}}\leq b^{i}_{o}$) as